Circular No. 8739

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V4332 SAGITTARII D. K. Lynch, R. W. Russell, and R. Ford, The Aerospace Corporation;

H.B. Hammel, Space Science Institute (SSI); and M. L. Sitko, University of Cincinnati and SSI, report 3- to 14-μm spectroscopy of V4332 Sgr (cf. IAUC 5944, 5945) on Aug. 7.35 UT using BASS at the Infrared Telescope Facility. V4332 Sgr showed a moderate infrared excess between 10 and 13 μm. The excess was nearly flat in units of W cm^{−2} μm^{−1} vs. wavelength, with a slight increasing brightness trend toward longer wavelengths. Possible weak emission features were present at 11.1 and 12.7 μm. There was a quasicontinuum between 3 and 9 μm that was too narrow to be fitted by a black-body function. Derived magnitudes for V4332 Sgr were L = 6.7 ± 0.3, M = 5.7 ± 0.2, N [narrow, 10.3 μm] = 5.8 ± 0.5.

COMETS C/2006 L6–L8, C/2006 M5–M9, C/2006 N1–N3 (SOHO) Additional Kreutz sungrazing comets have been found on SOHO web-

site images (cf. IAUC 8738) — all very faint except where noted below. C/2006 L6, C/2006 M5, C/2006 M8, and C/2006 M9 were slightly diffuse (and generally small), while C/2006 L7, C/2006 L8, and C/2006 M7 were diffuse. C/2006 M9 was also found by T. Chen. C/2006 L5–L8 and C/2006 M5–M7 peaked at mag ∼ 7.5 or fainter. C/2006 N1 was tiny, faint, and stellar in appearance in C3 images; in C2 images, it was diffuse and very faint with a hint of elongation. C/2006 N2 was small and stellar in appearance, peaking at mag ∼ 6.5. C/2006 N3 appeared stellar in C3 images, peaking at mag ≈ 5; in C2 images, it was rather condensed with no tail, fading rapidly.

Comet 2006 UT α_{2000} δ_{2000} Inst. F MPEC
C/2006 L6 June 8.833 5^{h}04.^{m}9 +21^{o}00′ C2  GS 2006-O21
C/2006 L7  10.450 5 10.9 +21 07 C2  TH 2006-O21
C/2006 L8  14.617 5 27.1 +21 24 C2  HS 2006-O21
C/2006 M5  16.521 5 34.6 +21 30 C2  TC 2006-O21
C/2006 M6  19.064 5 44.4 +21 38 C2  TH 2006-O21
C/2006 M7  22.438 5 57.2 +21 40 C2  HS 2006-O21
C/2006 M8  25.064 6 07.5 +21 47 C2  HS 2006-O62
C/2006 M9  27.163 6 15.7 +21 44 C2  HS 2006-O62
C/2006 N1 July 11.821 7 12.0 +20 10 C3/2 SF 2006-O62
C/2006 N2 13.971 7 17.4 +19 55 C3 HS 2006-O62
C/2006 N3  14.363 7 19.5 +17 46 C3/2 WX 2006-O62
2006 August 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green